# M2SE: A Multistage Multitask Instruction Tuning Strategy for Unified Sentiment and Emotion Analysis

### v1

@@ II-B Multitask Learning (MTL)
Multitask Learning (MTL) is a machine learning paradigm that enables models to simultaneously learn multiple related tasks, thereby allowing knowledge from one task to be transferred to others [41]. In affective computing, where multimodal inputs (e.g., text, audio, and visual frames) carry diverse and distinct information, MTL frameworks aim to leverage the synergies between tasks for improved performance. For instance, the MTMM-ES framework [42] jointly performs sentiment and emotion analysis, showing improved results over single-task models. UniMSE [11] unifies MSA and ER within a single framework. However, it relies solely on labeled data to establish the relationship between sentiment and emotion, without fully utilizing multimodal connections. M2Seq2Seq [43] proposes a multimodal multitask learning model based on the encoder-decoder architecture. Recently, Emotion-LLaMA [39] has shown promising results in ERC by incorporating multitask learning, including emotion cause inference. However, despite its success in ERC, Emotion-LLaMA still struggles with MSA tasks and fails to effectively synthesize contextual information for emotion inference.
@@ IV-D Ablation Evaluation
Model CMU-MOSEI MELD ECF2.0 Acc2 Acc W.F1 F1 W.F1 EmoVerse(Z.S.) 82.09 / 83.87 39.46 40.59 33.24 33.35 EmoVerse(MSA) 83.96 / 88.16 48.12 31.27 41.67 41.14 EmoVerse(ER) 78.99 / 80.17 66.21 65.71 40.61 38.23 EmoVerse(EPCE) 83.27 / 85.85 32.38 33.43 72.69 72.59 EmoVerse(M2SE) 85.51 / 87.80 67.78 66.74 72.30 72.21

### v2

@@ II-B Multitask Learning (MTL)
Multitask Learning (MTL) is a machine learning paradigm that enables models to simultaneously learn multiple related tasks, thereby allowing knowledge from one task to be transferred to others [41]. In affective computing, where multimodal inputs (e.g., text, audio, and visual frames) carry diverse and distinct information, MTL frameworks aim to leverage the synergies between tasks for improved performance. For instance, the MTMM-ES framework [42] jointly performs sentiment and emotion analysis, showing improved results over single-task models. UniMSE [11] unifies MSA and ER within a single framework. However, it relies solely on labeled data to establish the relationship between sentiment and emotion, without fully utilizing multimodal connections. M2Seq2Seq [43] proposes a multimodal multitask learning model based on the encoder-decoder architecture. Recently, Emotion-LLaMA [39] has shown promising results in MER by incorporating multitask learning, including emotion reason inference. However, despite its success in MER, Emotion-LLaMA still struggles with MSA tasks and fails to effectively synthesize contextual information for emotion inference.
@@ IV-D Ablation Evaluation
Model CMU-MOSEI MELD ECF2.0 Acc2 Acc W.F1 F1 W.F1 EmoVerse(Z.S.) 82.09 / 83.87 39.46 40.59 33.24 33.35 EmoVerse(MSA) 83.96 / 88.16 48.12 31.27 41.67 41.14 EmoVerse(ER) 78.99 / 80.17 66.21 65.71 40.61 38.23 EmoVerse(EPCE) 83.27 / 85.85 32.38 33.43 72.69 72.59 EmoVerse(M2SE) 85.51 / 87.80 67.28 66.58 72.30 72.21